Can you please help me with this one? I think it's linear. ** A rectangle has a length 5 inches longer than its width. If both the length and the width of the rectangle are increased by 3 inches, the area of the rectangle is increased by 36 square inches. Find the original dimensions of the rectangle.
Answer by venugopalramana(3286)

Can you please help me with this one? I think it's linear. ** A rectangle has a length 5 inches longer than its width.
LET WIDTH =W
LENGTH=W+5
AREA=L*B=W(W+5)=W^2+5W
If both the length and the width of the rectangle are increased by 3 inches,
L NEW =L+3=W+5+3=W+8
W NEW =W+3
AREA NEW =(W+8)(W+3)= W^2+8W+3W+24=W^2+11W+24
the area of the rectangle is increased by 36 square inches.
W^2+11W+24-W^2-5W=36
6W=36-24=12
W=12/6=2
L=W+5=2+5=7
Find the original dimensions of the rectangle.